# Unleashing the Potential of Vision-Language Pre-Training for 3D Zero-Shot Lesion Segmentation via Mask-Attribute Alignment

### v1

@@ Abstract
Recent advancements in medical vision-language pre-training models have driven significant progress in zero-shot disease recognition. However, transferring image-level knowledge to pixel-level tasks, such as lesion segmentation in 3D CT scans, remains a critical challenge. Due to the complexity and variability of pathological visual characteristics, existing methods struggle to align fine-grained lesion features not encountered during training with disease-related textual representations. In this paper, we present Malenia, a novel multi-scale lesion-level mask-attribute alignment framework, specifically designed for 3D zero-shot lesion segmentation. Malenia improves the compatibility between mask representations and their associated elemental attributes, explicitly linking the visual features of unseen lesions with the extensible knowledge learned from previously seen ones. Furthermore, we design a Cross-Modal Knowledge Injection module to enhance both visual and textual features with mutually beneficial information, effectively guiding the generation of segmentation results. Comprehensive experiments across three datasets and 12 lesion categories validate the superior performance of Malenia. Codes will be publicly available.
11footnotetext: Equal contribution. 🖂Corresponding authors. This work was supported by Shanghai Artificial Intelligence Laboratory.
@@ 1 Introduction
Motivated by the aforementioned limitations, we introduce Malenia, a novel multi-scale lesion-level mask-attribute alignment framework for superior zero-shot lesion segmentation. Malenia first leverages multi-scale mask representations with inherent boundary information to capture diverse lesion regions, then matches fine-grained visual features of lesions with text embeddings, effectively bridging the gap between the contrastive pre-training task and the per-pixel dense prediction task. To learn extensible representations that are robust to the out-of-distribution visual characteristics of unseen lesions, we incorporate domain knowledge from human experts to structure textual reports into descriptions of various elemental disease visual attributes, such as shape, location, density, and specific patterns related to disease manifestations. Despite the significant variability among lesions, these fundamental attributes are shared across different diseases and are often represented similarly in images. By aligning mask representations of lesions with their corresponding visual attributes, the model mimics the decision-making process of human experts, explicitly linking the visual features of unseen diseases to the intrinsic attributes learned from seen lesions. This alignment significantly enhances the model’s ability to recognize and segment various unseen lesions. Furthermore, we propose a novel Cross-Modal Knowledge Injection (CMKI) module in Malenia, inspired by the observation that visual and textual embeddings, after feature alignment, are complementary and can mutually reinforce each other. Unlike existing language-driven segmentation methods (Liu et al., 2023; Jiang et al., 2024), which keep text embeddings fixed, the CMKI module updates both mask and attribute embeddings to facilitate fine-grained multi-modal feature fusion. We leverage both enhanced mask and attribute embeddings to generate predictions by matching query features with image features, and then ensemble these predictions to produce the final segmentation results, demonstrating improved performance for both seen and unseen lesions. To thoroughly validate the effectiveness of Malenia, we evaluate its segmentation performance on both seen and unseen lesions using the MSD (Antonelli et al., 2022), KiTS23 (Heller et al., 2023), and a curated real-world in-house dataset. Malenia consistently outperforms state-of-the-art methods across 12 lesion categories. Our contributions can be summarized as follows:
  • We present Malenia, a novel multi-scale lesion-level mask-attribute alignment framework that captures extensible multi-modal representations for significantly improved zero-shot lesion segmentation by effectively matching the fine-grained visual appearances of new diseases with the textual representations of various fundamental pathological visual attributes.
@@ 2 Related Works
Zero-Shot 3D Medical Image Segmentation. Motivated by the impressive zero-shot performance of SAM (Kirillov et al., 2023) and SAM 2 (Ravi et al., 2024) in natural images, numerous studies have evaluated their application in medical image segmentation (Wald et al., 2023; Zhang et al., 2023; Huang et al., 2024; Yamagishi et al., 2024). However, due to the significant domain gap, directly applying SAM or SAM 2 to medical images typically results in unsatisfactory performance. Consequently, recent works (Ma et al., 2024; Guo et al., 2024; Shen et al., 2024; Zhu et al., 2024; Shaharabany & Wolf, 2024) explore effective adaptations of prompted SAM or SAM 2 on medical datasets. Nevertheless, these SAM adaptations require prompts (points, bounding boxes, or masks) sampled from ground truth during testing, which demands significant expertise and is often impractical in real-world clinical scenarios. As a result, prompt-free SAM adaptation methods (Hu et al., 2023; Zhang & Liu, 2023; Cheng et al., 2024; Aleem et al., 2024) have also been proposed. To the best of our knowledge, although the current SAM-based 3D medical image segmentation methods demonstrate promising zero-shot performance in segmenting certain organs in CT scans, they have not been evaluated or proven effective when confronted with unseen lesions that have less defined structures. Apart from SAM-based models, a self-prompted method, ZePT (Jiang et al., 2024), achieves competitive zero-shot tumor segmentation performance by matching class-agnostic mask proposals with text descriptions of general medical knowledge. However, this method does not leverage patient-specific information from reports and overlooks the multi-aspect visual attributes shared across different diseases, leading to compromised zero-shot performance. Our approach goes a step further, achieving significantly superior zero-shot lesion segmentation performance without the need for complex visual prompts.
@@ 3 Method
Fig. 1 illustrates the pipeline of Malenia. It is built upon the recent mask-based segmentation method Mask2Former (Cheng et al., 2022), which has achieved promising performance in the field of medical image segmentation (Yuan et al., 2023; Chen et al., 2023; Jiang et al., 2024). Below, we introduce (i) our novel multi-scale mask-attribute alignment strategy (Sec.3.1) that effectively aligns lesion representations with text embeddings of fundamental disease attributes; (ii) the Cross-Modal Knowledge Injection module (Sec.3.2) for multi-modal feature representation enhancement; and (iii) the overall training objectives and inference details (Sec. 3.3). We provide the preliminaries, discussing Mask2Former and the problem formulation of zero-shot lesion segmentation, as prior information required for our work in Appendix A.
Figure 1: Overview of the training and inference framework of Malenia. The key contributions of our work are two simple but effective designs: the Multi-scale Fine-Grained Mask-Attribute Alignment and the Cross-Modal Knowledge Injection module. Incorporating these designs empowers our vision-language pre-training framework, resulting in a strong zero-shot segmentation model.
@@ 3.1 Multi-Scale Mask-Text Contrastive Learning
$\centering\displaystyle M^{(i)}_{cross}=\text{CrossAttn}(q^{(i)},k^{(i)},v^{(i)}),\quad M^{(i+1)}=\text{SelfAttn}(M^{(i)}_{cross})\@add@centering$ (1)
@@
(1) Utilization of Multi-Scale Features. Existing methods (Jiang et al., 2024; Lin et al., 2024) overlook the advantage of leveraging multi-scale visual features during cross-modal alignment. In contrast, we match the hierarchical mask token embeddings from different Transformer decoder blocks with text features. This approach enables the model to accurately capture and segment across a range of mask sizes. Specifically, the mask tokens interact with image features whose dimensions are set as $(h^{i},w^{i},D^{i})=(H/32,W/32,D/32),(H/16,W/16,D/16),(H/8,W/8,D/8)$ for blcoks $i=1,2,3$, respectively. This variation in feature resolution across blocks ensures mask-text alignment at different scales, which is crucial for segmenting classes with large size variations, such as tumors.
@@ 3.2 Cross-Modal Knowledge Injection Module
$Mask_{m}=\frac{Q_{m}K^{T}}{\sqrt{C_{k}}}\in\mathbb{R}^{H\times W\times D\times N},\quad Mask_{t}=\frac{Q_{t}K^{T}}{\sqrt{C_{k}}}\in\mathbb{R}^{H\times W\times D\times N}$ (5)
Here, $Mask_{m}$ and $Mask_{t}$ refer to the two-branch output mask predictions derived from mask tokens and text embeddings, respectively. $\sqrt{C_{k}}$ is the dimension of the keys as a scaling factor. We ensemble these mask predictions as $Mask=MLP(\beta_{1}Mask_{m}+\beta_{2}Mask_{t})$. Here $\beta_{1}$ and $\beta_{2}$ are weighting factors for the vision and language branches, respectively, set to $0.5$ by default. The final segmentation results are obtained by applying the Argmax operation along the channel dimension of the $Mask$.
@@ 3.3 Training Objectives and Inference
Inference: Fig. 1 (b) illustrates the model testing flow. To enable convenient and flexible inference, we store the extracted text features of all visual attribute descriptions after training, eliminating the need for the text encoder during testing and thereby reducing computational overhead. The $N$ mask tokens first interact with image features from the 3D image decoder to identify all potential lesion regions. Next, by calculating the similarity between mask tokens and stored text features, we match the mask tokens with their corresponding attribute embeddings (selecting the highest similarity score for each attribute aspect). These matched visual and textual tokens are then fed into the CMKI module to generate the final segmentation results. Based on the matched eight different visual attributes, which can essentially be viewed as diagnostic criteria axes, we obtain the class information for each predicted lesion mask by referencing a clinical knowledge table that records the correspondence between each disease and its eight visual attributes.
@@ 4.1 Experiment Setup
Datasets. We utilize annotated datasets encompassing lesions from $12$ diseases and $6$ organs, sourced from both public and private datasets. Specifically, we consider the MSD dataset (Antonelli et al., 2022) and KiTS23 (Heller et al., 2023) dataset. We also collect a private dataset that includes four lesion types: liver cyst, gallbladder tumor, gallstone, and kidney stone. For training, we use the following datasets: 1) colon tumor, lung tumor, liver tumor, and pancreas tumor from the MSD dataset; 2) kidney cyst from the KiTS23 dataset; 3) gallstone from our private dataset. For zero-shot segmentation testing, we adopt: 1) hepatic vessel tumor and pancreas cyst from the MSD dataset; 2) kidney tumor from the KiTS23 dataset; and 3) liver cyst, gallbladder tumor and kidney stone from our private dataset. Details of all datasets, the annotation process for disease attribute descriptions, and preprocessing are provided in Appendix B.
Evaluation Metrics. We adopt standard segmentation metrics, including the Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD). Additionally, we report the the computational efficiency evaluation, which are detailed in Appendix E.
@@ 4.2 Main Results
Segmentation performance on seen lesions. We also evaluate the segmentation performance of Malenia on seen lesions in the fully-supervised setting. We compare Malenia with SOTA medical image segmentation methods, including TransUNet (Chen et al., 2021), nnUNet (Isensee et al., 2021), Swin UNETR (Tang et al., 2022), and the Universal Model (Liu et al., 2023). As shown in Table 2, Malenia outperforms competing baselines on seen lesions, achieving an average improvement of $1.92\%$ in DSC for tumors on the MSD dataset (Antonelli et al., 2022) and $1.35\%$ in DSC for kidney cysts on KiTS23 (Heller et al., 2023). Most segmentation baselines (e.g., TransUNet, nnUNet, and Swin UNETR) focus solely on visual features and overlook the semantic relationships between different lesions. Although the Universal Model incorporates text embeddings of category names to help learn correlations between anatomical regions, it still lacks fine-grained information from both the vision and language domains. In contrast, Malenia outperforms these methods by aligning fine-grained lesion-level semantics with comprehensive, patient-specific textual features and enhancing mask representations through cross-modal knowledge injection. This improvement demonstrates that our novel strategies also enhance the segmentation of seen lesion categories.
@@
TransUNet* 44.78±16.21 54.14±15.67 38.85±10.25 54.72±11.59 60.05±5.29 72.88±5.98 53.39±6.76 55.94±7.35 48.43±14.04 52.32±15.62
nnUNet* 47.02±15.85 57.36±14.33 37.97±10.54 53.98±11.86 61.33±5.01 73.27±5.44 54.44±6.30 56.48±7.17 48.76±13.82 52.96±15.19
Swin UNETR* 46.87±16.02 55.28±15.52 38.72±10.33 54.01±11.67 62.37±4.88 74.75±5.09 55.65±6.18 57.98±6.54 48.06±14.26 52.11±16.05
Universal Model* 51.02±14.62 60.93±13.36 42.40±9.54 58.54±10.79 64.25±3.94 77.06±4.21 54.51±6.27 56.68±7.11 50.25±12.24 54.17±13.53
Malenia 53.55±13.49 62.41±12.81 43.30±9.29 59.63±10.55 65.18±3.74 78.95±4.03 57.84±5.53 59.97±5.90 51.60±11.84 55.41±12.99

### v2

@@ Abstract
Recent advancements in medical vision-language pre-training models have driven significant progress in zero-shot disease recognition. However, transferring image-level knowledge to pixel-level tasks, such as lesion segmentation in 3D CT scans, remains a critical challenge. Due to the complexity and variability of pathological visual characteristics, existing methods struggle to align fine-grained lesion features not encountered during training with disease-related textual representations. In this paper, we present Malenia, a novel multi-scale lesion-level mask-attribute alignment framework, specifically designed for 3D zero-shot lesion segmentation. Malenia improves the compatibility between mask representations and their associated elemental attributes, explicitly linking the visual features of unseen lesions with the extensible knowledge learned from previously seen ones. Furthermore, we design a Cross-Modal Knowledge Injection module to enhance both visual and textual features with mutually beneficial information, effectively guiding the generation of segmentation results. Comprehensive experiments across three datasets and 12 lesion categories validate the superior performance of Malenia.
11footnotetext: Equal contribution. 🖂Corresponding authors. This work was supported by Shanghai Artificial Intelligence Laboratory. Codes are available at https://github.com/Yankai96/Malenia.
@@ 1 Introduction
Motivated by the aforementioned limitations, we introduce Malenia, a novel multi-scale lesion-level mask-attribute alignment framework for superior zero-shot lesion segmentation. Malenia first leverages multi-scale mask representations with inherent boundary information to capture diverse lesion regions, then matches fine-grained visual features of lesions with text embeddings, effectively bridging the gap between the contrastive pre-training task and the per-pixel dense prediction task. To learn extensible representations that are robust to the out-of-distribution visual characteristics of unseen lesions, we incorporate domain knowledge from human experts to structure textual reports into descriptions of various elemental disease visual attributes (e.g., shape, intensity, location). Despite the significant variability among lesions, these fundamental attributes are shared across different diseases and are often represented similarly in images. By aligning mask representations of lesions with their corresponding visual attributes, the model mimics the decision-making process of human experts, explicitly linking the visual features of unseen diseases to the intrinsic attributes learned from seen lesions. Furthermore, we propose a novel Cross-Modal Knowledge Injection (CMKI) module in Malenia, inspired by the observation that visual and textual embeddings, after feature alignment, are complementary and can mutually reinforce each other. The CMKI module updates both mask and attribute embeddings to facilitate fine-grained multi-modal feature fusion. We leverage both enhanced mask and attribute embeddings to generate predictions by matching query features with image features, and then ensemble these predictions to produce the final segmentation results, demonstrating improved performance for both seen and unseen lesions. To thoroughly validate the effectiveness of Malenia, we evaluate its segmentation performance on both seen and unseen lesions using the MSD (Antonelli et al., 2022), KiTS23 (Heller et al., 2023), and a curated real-world in-house dataset. Malenia consistently outperforms state-of-the-art methods across 12 lesion categories. Our contributions can be summarized as follows:
  • We present Malenia, a novel multi-scale lesion-level mask-attribute alignment framework that captures extensible multi-modal representations for significantly improved zero-shot lesion segmentation by effectively matching the fine-grained visual appearances of new diseases with the textual representations of various fundamental pathological attributes.
@@ 2 Related Works
Zero-Shot 3D Medical Image Segmentation. Motivated by the impressive zero-shot performance of SAM (Kirillov et al., 2023) and SAM 2 (Ravi et al., 2024) in natural images, numerous studies have evaluated their application in medical image segmentation (Wald et al., 2023; Zhang et al., 2023; Huang et al., 2024b; Yamagishi et al., 2024) and explored effective adaptations of SAM or SAM 2 on medical datasets (Ma et al., 2024; Guo et al., 2024; Shen et al., 2024; Zhu et al., 2024; Shaharabany & Wolf, 2024). Nevertheless, these SAM-based medical image segmentation methods require prompts (points, bounding boxes, or masks) sampled from ground truth during testing, which demands significant expertise and is often impractical in real-world clinical scenarios. As a result, prompt-free SAM adaptation methods (Hu et al., 2023; Zhang & Liu, 2023; Cheng et al., 2024; Aleem et al., 2024) have also been proposed. To the best of our knowledge, although the current SAM-based 3D medical image segmentation methods demonstrate promising zero-shot performance in segmenting certain organs in CT scans, they have not been evaluated or proven effective when confronted with unseen lesions that have less defined structures. Apart from SAM-based models, a self-prompted method, ZePT (Jiang et al., 2024), achieves competitive zero-shot tumor segmentation performance by matching class-agnostic mask proposals with text descriptions of general medical knowledge. However, it overlooks patient-specific information from reports, leading to compromised zero-shot performance. Our approach goes a step further, achieving significantly superior zero-shot lesion segmentation performance without the need for complex visual prompts.
Language-Guided Medical Image Segmentation. Recently, several methods also incorporate vision and language information to enhance medical image segmentation (Li et al., 2023; Liu et al., 2023; Huang et al., 2024a). Notably, LViT (Li et al., 2023) leverages medical texts to compensate for quality deficiencies in image data and guide the generation of pseudo labels in semi-supervised settings. RecLMIS (Huang et al., 2024a) introduces conditioned reconstruction to explicitly capture cross-modal interactions between medical images and text, leading to enhanced performance. However, these methods primarily focus on fully-supervised and semi-supervised settings, aiming to enhance segmentation performance for seen categories by combining visual and textual features. In contrast, Malenia targets the zero-shot setting, aiming to develop generalized and robust textual features, such as disease attributes, along with a novel vision-language alignment strategy to enhance segmentation performance for unseen lesions.
@@ 3 Method
Fig. 1 illustrates the pipeline of Malenia. It is built upon the recent mask-based segmentation backbone Mask2Former (Cheng et al., 2022). Below, we introduce (i) our novel multi-scale mask-attribute alignment strategy (Sec.3.1) that effectively aligns lesion representations with text embeddings of fundamental disease attributes; (ii) the Cross-Modal Knowledge Injection module (Sec.3.2) for multi-modal feature representation enhancement; and (iii) the overall training objectives and inference details (Sec. 3.3). We provide the preliminaries, discussing Mask2Former and the problem formulation of zero-shot lesion segmentation, as prior information required for our work in Appendix A.
Figure 1: Overview of Malenia. The key contributions of our work are two simple but effective designs: the Multi-scale Fine-Grained Mask-Attribute Alignment and the CMKI module.
@@ 3.1 Multi-Scale Mask-Text Contrastive Learning
$M^{(i)}_{cross}=\text{CrossAttn}(q^{(i)},k^{(i)},v^{(i)}),\quad M^{(i)}=\text{SelfAttn}(M^{(i)}_{cross})$ (1)
@@
(1) Utilization of Multi-Scale Features. Existing methods (Jiang et al., 2024; Lin et al., 2024) overlook the advantage of leveraging multi-scale visual features during cross-modal alignment. In contrast, we match the hierarchical mask token embeddings from different Transformer decoder blocks with text features. Specifically, the mask tokens interact with image features whose dimensions are set as $(h^{i},w^{i},D^{i})=(H/32,W/32,D/32),(H/16,W/16,D/16),(H/8,W/8,D/8)$ for blcoks $i=1,2,3$, respectively. This variation in feature resolution across blocks ensures mask-text alignment at different scales, which is crucial for segmenting classes with large size variations, such as tumors.
@@ 3.2 Cross-Modal Knowledge Injection Module
$Mask_{m}=\frac{Q_{m}K^{T}}{\sqrt{C}}\in\mathbb{R}^{H\times W\times D\times N},\quad Mask_{t}=\frac{Q_{t}K^{T}}{\sqrt{C}}\in\mathbb{R}^{H\times W\times D\times N}$ (5)
Here, $Mask_{m}$ and $Mask_{t}$ refer to the two-branch output mask predictions derived from mask tokens and text embeddings, respectively. $\sqrt{C}$ is the dimension of the keys as a scaling factor. We ensemble these mask predictions as $Mask=MLP(\beta_{1}Mask_{m}+\beta_{2}Mask_{t})$. Here $\beta_{1}$ and $\beta_{2}$ are weighting factors for the vision and language branches, respectively, set to $0.5$ by default. The final segmentation results are obtained by applying the Argmax operation along the channel dimension of the $Mask$.
Figure 2: Overview of the inference process of Malenia. (1) Step-I: Image Partitioning via Mask Tokens. Test CT images are divided into regions, each represented by mask tokens. (2) Step-II: Mask-attribute matching. Each mask token is associated with stored attribute embeddings. (3) Step-III: Cross-modal fusion and mask prediction. Information from mask tokens and text embeddings is fused to generate segmentation masks. (4) Step-IV: Disease identification via attribute-querying. The Clinical Knowledge Table links the predicted attributes to specific disease categories for precise diagnosis.
@@ 3.3 Training Objectives and Inference
Inference: Fig. 2 illustrates the model testing flow. To enable convenient and flexible inference, we store the text features of all visual attribute descriptions after training, eliminating the need for the text encoder during testing. The inference process can be divided into four steps. Step-I: Mask tokens attend to different regions among the images, partitioning the image into different regions. In this process, different lesion regions in the image are captured by specific mask tokens. Step-II: We compute the similarity between the $N$ mask tokens and the stored attribute embeddings, after which the mask tokens are matched to the text embeddings based on these similarity scores. Step-III: The matched mask tokens and text tokens are fed into the CMKI module for further feature fusion. Then, the information from mask tokens and text embeddings are combined to produce the final mask prediction. Step-IV: The text attributes of each mask are used to identify the specific category. This is achieved by querying a Clinical Knowledge Table that records the relationships between each disease and its attributes. The Clinical Knowledge Table are given in Table 11.
@@ 4.1 Experiment Setup
Datasets. We utilize annotated datasets encompassing lesions from $12$ diseases and $6$ organs, sourced from both public and private datasets. (1) Public datasets: We consider the MSD dataset (Antonelli et al., 2022) and KiTS23 (Heller et al., 2023) dataset. Two senior radiologists supplemented the annotated lesions in MSD and KiTS23 datasets with diagnostic reports and corresponding lesion descriptions of the eight visual attribute aspects. We also collect a private dataset***This retrospective study was approved by the ethics committee of the First Hospital of China Medical University; the informed consent requirement was waived. All diagnostic reports in the private dataset used in this study have been fully anonymized, with all patient-identifiable information and private details removed. that includes four lesion types: liver cyst, gallbladder tumor, gallstone, and kidney stone. For training, we use the following datasets: 1) colon tumor, lung tumor, liver tumor, and pancreas tumor from the MSD dataset; 2) kidney cyst from the KiTS23 dataset; 3) gallstone from our private dataset. For zero-shot testing, we adopt: 1) hepatic vessel tumor and pancreas cyst from the MSD dataset; 2) kidney tumor from the KiTS23 dataset; and 3) liver cyst, gallbladder tumor and kidney stone from our private dataset. Details of all datasets, the annotation process for disease attribute descriptions, and preprocessing are provided in Appendix B.
Evaluation Metrics. We adopt standard segmentation metrics, including the Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD). Additionally, we report the the computational efficiency evaluation, which are detailed in Appendix G.
@@ 4.2 Main Results
Segmentation performance on seen lesions. We also evaluate the segmentation performance of Malenia on seen lesions in the fully-supervised setting. We compare Malenia with SOTA medical image segmentation methods, including TransUNet (Chen et al., 2021), nnUNet (Isensee et al., 2021), Swin UNETR (Tang et al., 2022), and the Universal Model (Liu et al., 2023). As shown in Table 2, Malenia outperforms competing baselines on seen lesions, achieving an average improvement of $1.46\%$ in DSC for tumors on the MSD dataset (Antonelli et al., 2022) and $1.35\%$ in DSC for kidney cysts on KiTS23 (Heller et al., 2023). Most segmentation baselines (e.g., TransUNet, nnUNet, and Swin UNETR) focus solely on visual features and overlook the semantic relationships between different lesions. Although the Universal Model incorporates text embeddings of category names to learn correlations between anatomical regions, it lacks fine-grained information from both the vision and language domains. In contrast, Malenia outperforms these methods by aligning fine-grained lesion-level semantics with comprehensive, patient-specific textual features and enhancing mask representations through cross-modal knowledge injection. This improvement demonstrates that our novel strategies also enhance the segmentation of seen lesion categories.
@@
TransUNet* 44.78±16.21 54.14±15.67 38.85±10.25 54.72±11.59 60.05±5.29 72.88±5.98 67.13±6.08 68.89±7.22 48.43±14.04 52.32±15.62
nnUNet* 47.02±15.85 57.36±14.33 37.97±10.54 53.98±11.86 61.33±5.01 73.27±5.44 69.50±5.61 71.39±6.55 48.76±13.82 52.96±15.19
Swin UNETR* 46.87±16.02 55.28±15.52 38.72±10.33 54.01±11.67 62.37±4.88 74.75±5.09 68.95±5.67 71.03±6.82 48.06±14.26 52.11±16.05
Universal Model* 51.02±14.62 60.93±13.36 42.40±9.54 58.54±10.79 64.25±3.94 77.06±4.21 67.27±5.71 69.33±6.95 50.25±12.24 54.17±13.53
Malenia 53.55±13.49 62.41±12.81 43.30±9.29 59.63±10.55 65.18±3.74 78.95±4.03 70.96±5.56 72.34±6.29 51.60±11.84 55.41±12.99